# Grounded GUI Understanding for Vision Based Spatial Intelligent Agent: Exemplified by Virtual Reality Apps

### v3

@@ Abstract.
In recent years, spatial computing Virtual Reality (VR) has emerged as a transformative technology, offering users immersive and interactive experiences across diversified virtual environments. Users can interact with VR apps through interactable GUI elements (IGEs) on the stereoscopic three-dimensional (3D) graphical user interface (GUI). The accurate recognition of these IGEs is instrumental, serving as the foundation of many software engineering tasks, including automated testing and effective GUI search. The most recent IGE detection approaches for 2D mobile apps typically train a supervised object detection model based on a large-scale manually-labeled GUI dataset, usually with a pre-defined set of clickable GUI element categories like buttons and spinners. Such approaches can hardly be applied to IGE detection in VR apps, due to a multitude of challenges including complexities posed by open-vocabulary and heterogeneous IGE categories, intricacies of context-sensitive interactability, and the necessities of precise spatial perception and visual-semantic alignment for accurate IGE detection results. Thus, it is necessary to embark on the IGE research tailored to VR apps.
In this paper, we propose the first zero-shot cOntext-sensitive inteRactable GUI ElemeNT dEtection framework for virtual Reality apps, named Orienter. By imitating human behaviors, Orienter observes and understands the semantic contexts of VR app scenes first, before performing the detection. The detection process is iterated within a feedback-directed validation and reflection loop. Specifically, Orienter contains three components, including (1) Semantic context comprehension for capturing the apps’ GUI context, (2) Reflection-directed IGE candidate detection for identifying and localizing valid GUI elements based on multi-perspective description guided IGE detection, as well as feedback-directed reflection, and (3) Context-sensitive interactability classification which integrates semantic contexts for interactability prediction. To evaluate our approach and facilitate follow-up research, we spend more than three months constructing the first benchmark dataset which contains 1,552 images from 100 industrial-setting apps on Steam, with 4,470 interactable annotations across 766 semantics categories. Extensive experiments on the dataset demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches (i.e., GPT-4o, YOLO v8, CenterNet2, Faster R-CNN, UIED, and Xianyu), surpassing their F1 Score by up to 3.7$\times$ and 121.4$\times$ (1.4$\times$ and 46.2$\times$ on average) in distinguishing the interactibility and semantics of the IGEs, respectively. Orienter is beneficial for boosting the performance of automatic testing by isolating the interactable action space from the whole space, regardless of the testing strategies employed. Experiments demonstrate that Orienter-guided testing covers 103.1% more IGEs with 125.7% more effective interactions than testing without action space isolation.
Virtual Reality, Spatial Intelligence, GUI Agent, Large Multimodal Model
ccs: Software and its engineering Software development techniquesccs: Computing methodologies Computer visionccs: Human-centered computing Virtual reality
@@ 1. Introduction
In recent years, Virtual Reality (VR) has emerged as a transformative technology, offering users immersive experiences across various virtual or virtual-real environments. This technological advancement has catalyzed the development of a myriad of VR apps, now numbering in tens of thousands (Li et al., 2023). These apps, which span a diverse array of domains, including skill training (web, 2021b), entertainment (web, 2021c, a), medical procedures (Khor et al., 2016), and military training (Bhagat et al., 2016), have attracted over 171 million users (Statista, 2022). This exponential growth underscores a critical need for robust software processes, including development, testing, and maintenance, particularly in high-reliability contexts such as healthcare and military training.
@@
In VR apps, users experience multimodal perceptions through various devices and interact via body movements and gestures. Among the multimodal perceptions, visual perception obtained via an app’s Graphical User Interface (GUI) offers the wealthiest information. GUIs in VR apps are often composed of three-dimensional (3D) GUI elements (e.g., images, text, widgets, etc.) or real-life objects and users interact with VR apps through Interactable GUI Elements (short as IGEs).
@@
Literature has shown that the accurate recognition of IGEs is a cornerstone for many software engineering tasks including automated testing (Xie et al., 2020; Ye et al., 2021; Chen et al., 2020) and effective GUI search (Reiss et al., 2018; Deka et al., 2017). For example, Ye et al. (Ye et al., 2021) report that 77% software experts believe that precise IGE detection can boost software testing efficiency by at least 50%. White et al. (White et al., 2019) demonstrate that IGE detection improves branch coverage by 42.5% compared to random testing.
@@
Deep learning (DL) based object detection approaches like Faster R-CNN (Ren et al., 2017), CenterNet (Zhou et al., 2021) and YOLO (Jocher et al., 2023), have demonstrated promising performance. With such advancements, recent work has taken steps to explore DL-based IGE detection approaches in mobile apps and desktop apps (White et al., 2019; Chen et al., 2020; Xie et al., 2020; Ye et al., 2021). For example, Wu et al. (Wu et al., 2023) explore advanced DL-based approaches, like CenterNet2, YOLOv3, and YOLOv5, for IGE detection on mobile apps. The approaches can be typically divided into three steps: (1) manually summarize a finite set of IGE categories such as buttons, spinners, switches, checkboxes, etc., (2) manually label a large dataset, and then (3) train an object detection model on the dataset. However, such training-based approaches heavily rely on large annotated datasets so that they can hardly be directly applied to IGE detection in VR apps. As shown in Figure 1(a) and Figure 1(b), annotating large-scale VR IGE datasets is challenging due to the unique interaction mechanism of VR apps. The data annotation process demands extensive efforts, compounded by VR’s complex hardware usage and exhaustive interactions.
@@
Recent advancements of pretrained large multimodal models (LMMs) (OpenAI, 2023; Gemini Team, 2023) have shown their remarkable abilities in downstream tasks including image/natural language comprehension, question answering and logical reasoning, without large training sets on specific problems. Such capabilities provide us with new opportunities to resolve the aforementioned limitation of lacking datasets. However, our preliminary experiments (detailed in  section 2.2) reveal that LMMs suffer from severe spatial semantic hallucinations, tending to generate contextually coherent but factually incorrect or unrealistic IGEs with inaccurate locations and amounts as responses. These are mainly stemmed from the following challenges of VR IGE detection problem:
@@
Challenge #1: Open-vocabulary and heterogeneous GUI element categories. LMM or DL approaches perform well on detecting a finite pre-defined set of IGE categories that frequently appear in their training sets and adhere to standard visual appearance patterns (e.g., buttons, sliders, and checkboxes) as well as unified interaction mechanisms (e.g., tapping and long-tapping). However, VR apps with diverse scenarios contain infinite open-vocabulary categories of IGEs that own different visual appearances and interaction mechanisms (heterogeneity), as shown in Figure 1(a). In the stereoscopic 3D VR scene, from time to time, IGEs occlude each other or are observed from non-front perspectives (e.g., Figure 3(a)), making the visual appearances of even the same IGE overly diverse and heterogeneous. Hence, it is hard for LMM or DL approaches to identify such open-vocabulary and heterogeneous VR GUI element categories.
@@
Challenge #2: Context-sensitive interactability. The interactability of GUI elements in VR apps is highly dependent on the semantic context of the VR app content and scenarios. As shown in Figure 1(c) and 1(d), respectively, an object like a tree might be interactable in one app (e.g., a tree-planting game where it needs to be picked up and planted) but non-interactable in another (e.g., as a mere background element in a fishing game). The IGE categories that LMM or DL approaches perform well when GUI elements with similar appearances behave mostly the same in distinct contexts, like buttons and checkboxes. We call them context-independent GUI elements. However, the interactability of GUI elements in VR apps is highly dependent on the app-specific and scenario-specific context, which means they cannot be uniformly determined by appearances but require context-sensitive reasoning instead.
@@
Challenge #3: Accurate IGE detection results require precise spatial perception and visual-semantic alignment. While capable of processing visual data and generating textual descriptions, LMM struggles with accurate spatial perception. Their ability to outline and label objects is limited, often resulting in imprecise localization, misaligned bounding boxes, etc. This spatial deficiency hinders their applicability in VR environments, where precise IGE localization is necessary.
@@
(a) IGE categories in a VR lab training app (web, 2024d) with diverse visual appearances and interaction mechanisms
@@
(b) Diversified IGE categories in Figure 1(a) with annotations, dataset annotation requires massive effort
(c) An interactable tree in a VR gardening (tree planting) game (web, 2024e)
(d) Several non-interactable trees in a VR fishing game (web, 2024g)
Figure 1. Examples to demonstrate the challenges of IGE detection for VR apps (web, 2024e, g, d)
@@
To tackle the challenges, we propose a cOntext-sensitive inteRactable GUI ElemeNT dEtection framework for virtual Reality apps (short as Orienter). Instead of directly identifying the IGEs in VR apps, Orienter is designed to imitate human behavior by comprehending the semantic context first before conducting the detection; and the detection process is iterated within a feedback-directed validation and reflection loop (i.e., two levels of “looking before leaping” in the title). Specifically, Orienter mainly contains three major modules. ❶ Semantic context comprehension: For identifying heterogeneous IGE categories in an unsupervised and context-sensitive way, we make Orienter to understand and combine the global context (app content overview, including genres, storylines, interaction mechanisms, etc.) and local context (the current VR scenario within the user’s field of view) of the VR scenes under analysis through in-context reasoning of LMM. This phase is the foundation of all the subsequent steps. ❷ Reflection-directed IGE candidate detection: For identifying and localizing valid IGEs, Orienter first mines multi-perspective characteristics of GUI elements leveraging LMM, and then uses these characteristic descriptions as guidance to detect corresponding elements. Based on the detection results, Orienter further performs feedback-directed reflection to validate and refine IGE candidates using a feedback loop. ❸ Context-sensitive interactabilitiy classification: To predict the interactibility of detected IGEs, Orienter incorporates the semantic context from the first module and performs chain-of-thought in-context classification.
For facilitating the evaluation, we spend more than three months constructing the first benchmark dataset for IGE detection in VR apps. The benchmark dataset consists of 1,552 images from 100 industrial-setting apps on Steam (web, 2023c), with 4,470 interactable annotations across 766 semantics categories. We evaluate Orienter in terms of predictions on interactability and semantics. Orienter outperforms all baselines (GPT-4o (web, 2024b), YOLO v8 (Jocher et al., 2023), CenterNet2 (Zhou et al., 2021), Faster R-CNN (Ren et al., 2017), UIED (Xie et al., 2020), Xianyu (Chen et al., 2020)) in almost all metrics, including precision, recall, and F1 score. Specifically, Orienter achieves improvements of up to 3.7$\times$ and 121.4$\times$ (1.4$\times$ and 46.2$\times$ on average) on F1 scores in distinguishing the interactability and semantics of IGEs, respectively, demonstrating Orienter’s strong effectiveness, efficiency, and reliability. To illustrate the usefulness of Orienter on downstream software engineering (SE) tasks, we further evaluate how effective it boosts the performance of automated testing. Orienter’s IGE detection results can be utilized to isolate the interactable action space from the whole input space, regardless of the testing strategies employed. Experiment results reveal that Orienter-guided testing covers 103.1% more IGEs with 125.7% more effective interactions than testing without action space isolation.
@@
  • To the best of our knowledge, we are the first to formulate the interactable GUI element (IGE) detection problem for stereoscopic 3D GUI. We leverage the power of large multimodal models in understanding and analyzing GUIs of VR software, and propose a zero-shot context-sensitive GUI element detection framework, Orienter. Orienter novelly undergoes semantic context comprehension, reflection-directed IGE candidate detection, and context-sensitive interactability classification to tackle the challenges of VR IGE detection.
@@
  • Extensive experiments have been conducted to verify the effectiveness of the proposed IGE detection framework. The results demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches, surpassing their F1 Score by up to 3.7$\times$ and 121.4$\times$ (1.4$\times$ and 46.2$\times$ on average) in distinguishing the interactibility and semantics of the IGEs, respectively. Experiments also illustrate that Orienter is beneficial for boosting GUI testing by covering 103.1% more IGEs with 125.7% more effective interactions than random testing.
@@
  • We construct the first dataset for VR IGE detection, including 1,552 images from 100 industrial-setting apps on Steam, with 4,470 interactable annotations across 766 semantics categories.
2. Preliminaries
2.1. Background of Virtual Reality (VR)
2.1.1. VR Devices
VR technologies provide immersive experiences that blend virtual elements with the physical world. These experiences are accessible via diverse devices including PCs and standalone systems, creating interactive environments that integrate real and virtual elements. Key VR devices include: (a) Head-mounted displays (HMDs), essential for VR experiences, offering immersive visuals and audio. (b) Two handheld controllers and gesture recognition systems enable users to interact with the virtual world via physical actions and gestures. (c) Tracking systems, comprising sensors and cameras, track user movement and orientation, crucial for a responsive VR experience.
2.1.2. VR Interaction
Interaction within VR environments can occur through various mechanisms: (a) Pressing buttons on controllers to manipulate GUI elements. (b) Touching or moving devices towards GUI elements, sometimes using extended virtual tools for remote interaction. (c) Gazing at GUI elements or employing eye-tracking technologies in HMDs to select or move elements. (d) Gestures and Movements to interact with the virtual environment through physical actions. These interaction modes highlight the evolving nature of human-computer interaction within VR, offering increasingly diverse and immersive experiences.
2.2. Motivating Examples
In this section, we present our motivational study to further demonstrate the technical challenges listed in section 1.
Two authors randomly sample 30 VR apps from the popular Steam VR app store (web, 2023c), select screenshots from them, and use the prompt template in Figure 2 to instruct the widely-used LMM GPT-4o (web, 2024b) to identify and locate IGEs in the images.
Figure 2. Propmt template for motivating examples.
As shown in figure 3(a), the mailbox is wrongly recognized as “Missile-like Object”. It is because the mailbox is viewed from an unusual angle, resulting in a special shape in the view and making it harder for the LMMs to identify semantics. This misunderstanding of GUI elements reveals the LMM’s inability to face the heterogeneous GUI elements in VR apps.
@@ 2.2. Motivating Examples
Figure 3. LMM gives deviated IGE detection results on VR scenes (web, 2024f, c) because of aforementioned challenges
For another example, figure 3(b), the bounding boxes in the image either cover only part of the object, locate objects mismatched with the label, or even mark the background. Although the LMM identifies most objects in the image, it fails to locate them well, revealing its insufficient spatial perception.
To summarize, the LMM alone cannot effectively detect IGEs in highly complex and varied VR apps. To boost the LMM on IGE detection in VR apps, it is necessary to provide the LMM with more context information and integrate the model with different techniques. Following this idea, we propose Orienter to imitate human behavior to enhance the LMM on detecting IGEs in VR apps.

### v4

# Grounded GUI Understanding for Vision-Based Spatial Intelligent Agent: Exemplified by Extended Reality Apps

@@ Abstract.
In recent years, Extended Reality (XR) has emerged as a transformative technology, offering users immersive and interactive experiences across diversified virtual or virtual-real environments. Users can interact with XR applications (apps) through interactable GUI elements (IGEs) on the stereoscopic three-dimensional (3D) graphical user interface (GUI). IGE constitutes the fundamental element of XR GUI, embodying rich semantic information. The accurate recognition and precise understanding of these IGEs is instrumental, serving as the foundation of GUI grounding, which can facilitates downstream tasks, including automated XR testing. A straightforward XR test generator can interact randomly within the app’s 3D environment, making it trapped in uninteractable space and resulting in an ineffective and inefficient testing process. In contrast, a more intelligent test generator, informed by the accurate locations and semantics of IGEs, can make wiser decisions on interaction targets and orders, forming test sequences that cover more functionalities faster. The most recent IGE detection approaches in SE are designed for 2D mobile apps and typically train a supervised object detection model based on a large-scale manually-labeled GUI dataset, usually with a pre-defined set of clickable GUI element categories like buttons and spinners. Such approaches can hardly be applied to IGE detection in XR apps, due to a multitude of challenges including complexities posed by open-vocabulary and heterogeneous IGE categories, intricacies of context-sensitive interactability, and the necessities of precise spatial perception and visual-semantic alignment for accurate IGE detection results. Thus, it is necessary to embark on the IGE research tailored to XR apps.
[…]
@@ 1. Introduction
In recent years, Extended Reality (XR) has emerged as a transformative technology, offering users immersive experiences across various virtual or virtual-real environments. This technological advancement has catalyzed the development of a myriad of XR apps  (Li et al., 2023). These apps, which span a diverse array of domains, including skill training (web, 2021b), entertainment (web, 2021c, a), medical procedures (Khor et al., 2016), and military training (Bhagat et al., 2016), have attracted over 171 million users (Statista, 2022). This exponential growth underscores a critical need for robust software processes, including development, testing, and maintenance, particularly in high-reliability contexts such as healthcare and military training.
@@
In XR apps, users experience multimodal perceptions through various devices and interact via body movements and gestures. Among the multimodal perceptions, visual perception obtained via an app’s Graphical User Interface (GUI) offers the wealthiest information. GUIs in XR apps are often composed of three-dimensional (3D) GUI elements (e.g., images, text, widgets, etc.) or real-life objects and users interact with XR apps through Interactable GUI Elements (short as IGEs).
@@
As an important part of GUI grounding to facilitate downstream tasks, literature has shown that the accurate recognition of IGEs is a cornerstone for many software engineering tasks including automated testing (Xie et al., 2020; Ye et al., 2021; Chen et al., 2020) and GUI code generation (Nguyen and Csallner, 2015; Moran et al., 2020; Chen et al., 2018). For example, Ye et al. (Ye et al., 2021) report that 77% software experts believe that precise IGE detection can boost software testing efficiency by at least 50%. White et al. (White et al., 2019) demonstrate that IGE detection improves branch coverage by 42.5% compared to random testing.
@@
Deep learning (DL) based object detection approaches (Ren et al., 2017; Zhou et al., 2021; Jocher et al., 2023) have demonstrated promising performance. With such advancements, recent work has taken steps to explore DL-based IGE detection approaches in mobile apps and desktop apps (White et al., 2019; Chen et al., 2020; Xie et al., 2020; Ye et al., 2021). For example, Wu et al. (Wu et al., 2023) explore advanced DL-based approaches, like CenterNet2, YOLOv3, and YOLOv5, for IGE detection on mobile apps. The approaches can be typically divided into three steps: (1) manually summarize a finite set of IGE categories such as buttons, input boxes, etc., (2) manually label a large dataset, and then (3) train an object detection model on the dataset. However, such training-based approaches heavily rely on large annotated datasets so that they can hardly be directly applied to IGE detection in XR apps. As shown in Figure 1(a) and Figure 1(b), annotating large-scale XR IGE datasets is challenging due to the unique interaction mechanism of XR apps. The data annotation process demands extensive efforts, compounded by XR’s complex hardware usage and exhaustive interactions.
@@
Recent advancements of pretrained large multimodal models (LMMs) (OpenAI, 2023; Gemini Team, 2023) have shown their remarkable abilities in downstream tasks including image/natural language comprehension, question answering and logical reasoning, without large training sets on specific problems. Such capabilities provide us with new opportunities to resolve the aforementioned limitation of lacking datasets. However, our preliminary experiments (detailed in  section 2.1) reveal that LMMs suffer from severe spatial semantic hallucinations, tending to generate contextually coherent but factually incorrect or unrealistic IGEs with inaccurate locations and amounts as responses. These are mainly stemmed from the following challenges of XR IGE detection problem:
@@
Challenge #1: Open-vocabulary and heterogeneous GUI element categories. LMM or DL approaches perform well on detecting a finite pre-defined set of IGE categories that frequently appear in their training sets and adhere to standard visual appearance patterns (e.g., buttons, sliders, and checkboxes) as well as unified interaction mechanisms (e.g., tapping and long-tapping). However, XR apps with diverse scenarios contain infinite open-vocabulary categories of IGEs that own different visual appearances and interaction mechanisms (heterogeneity), as shown in Figure 1(a). In the stereoscopic 3D XR scene, IGEs sometimes occlude each other or are observed from non-front perspectives (e.g., Figure 3(a)), making the visual appearances of even the same IGE overly diverse and heterogeneous. Hence, it is hard for LMM or DL approaches to identify such open-vocabulary and heterogeneous XR GUI element categories.
@@
Challenge #2: Context-sensitive interactability. The interactability of GUI elements in XR apps is highly dependent on the semantic context of the XR app content and scenarios. As shown in Figure 1(c) and 1(d), respectively, an object like a tree might be interactable in one app (e.g., a tree-planting game where it needs to be picked up and planted) but non-interactable in another (e.g., as a mere background element in a fishing game). The IGE categories that LMM or DL approaches perform well when GUI elements with similar appearances behave mostly the same in distinct contexts, like buttons and checkboxes. However, the interactability of GUI elements in XR apps is highly dependent on the app-specific and scenario-specific context, which means they cannot be uniformly determined by appearances but require context-sensitive reasoning instead.
@@
Challenge #3: Accurate IGE detection results require precise spatial perception and visual-semantic alignment. While capable of processing visual data and generating textual descriptions, LMM struggles with accurate spatial perception. Their ability to outline and label objects is limited, often resulting in imprecise localization, misaligned bounding boxes, etc. This spatial deficiency hinders their applicability in XR environments, where precise IGE localization is necessary.
@@
(a) IGE categories in lab training app (web, 2024e) with diverse visual appearances and interaction mechanisms
@@
(b) IGE in diversified categories in Figure 1(a) with annotations, dataset annotation requires massive effort
(c) An interactable tree in a XR gardening (tree planting) game (web, 2024f)
(d) Several non-interactable trees in a XR fishing game (web, 2024h)
Figure 1. Examples to demonstrate the challenges of IGE detection for XR apps (web, 2024f, h, e)
To tackle the challenges, we propose a cOntext-sensitive inteRactable GUI ElemeNT detection framework for Extended Reality apps (short as Orienter). Instead of directly identifying the IGEs in XR apps, Orienter is designed to imitate human behavior by comprehending the semantic context first before conducting the detection; and the detection process is iterated within a feedback-directed validation and reflection loop (i.e., two levels of “looking before leaping” in the title).
@@
Specifically, Orienter mainly contains three major modules. ❶ Semantic context comprehension: For identifying heterogeneous IGE categories in an unsupervised and context-sensitive way, we make Orienter to understand and combine the global context (app content overview, including genres, storylines, interaction mechanisms, etc.) and local context (the current XR scenario within the user’s field of view) of the XR scenes under analysis through in-context reasoning of LMM. This phase is the foundation of all the subsequent steps. ❷ Reflection-directed IGE candidate detection: For identifying and localizing valid IGEs, Orienter first mines multi-perspective characteristics of GUI elements leveraging LMM, and then uses these characteristic descriptions as guidance to detect corresponding elements. Based on the detection results, Orienter further performs feedback-directed reflection to validate and refine IGE candidates using a feedback loop. ❸ Context-sensitive interactabilitiy classification: To predict the interactibility of detected IGEs, Orienter incorporates the semantic context from the first module and performs chain-of-thought in-context classification.
[…]
@@
  • To the best of our knowledge, we are the first to formulate the interactable GUI element (IGE) detection problem for stereoscopic 3D GUI. We leverage the power of large multimodal models in understanding and analyzing GUIs of XR software, and propose a zero-shot context-sensitive GUI element detection framework, Orienter. Orienter novelly undergoes semantic context comprehension, reflection-directed IGE candidate detection, and context-sensitive interactability classification to tackle the challenges of XR IGE detection.
@@
  • We conduct extensive experiments to verify the effectiveness of the proposed IGE detection framework. The results demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches. Experiments show that Orienter can facilitate GUI testing.
@@
  • We construct the first dataset for XR IGE detection (web, 2025), including 1,552 images from 100 industrial-setting apps on Steam, with 4,470 interactable annotations across 766 semantics categories.
2. Background on Extended Reality (XR)
XR Devices. XR technologies offer immersive experiences that seamlessly blend virtual and physical worlds. These technologies can be accessed through various device types. VR typically requires either a PC, a standalone system, or a mobile phone for computational power; AR usually overlays digital information onto the physical world, often through smartphones or specialized AR glasses; and MR integrates the capabilities of both VR and AR, offering more complex and interactive experiences that merge real and virtual environments.
Typical XR devices include: (1) Head-mounted displays (HMDs), which encompass screens that display digital environments and often with audio capabilities. (2) Handheld controllers and gesture recognition systems, which allows users to navigate and manipulate the virtual environment through physical movements. (3) Tracking systems, which includes sensors and cameras, tracks user movements and spatial orientation, enabling a responsive and immersive experience in 3D space.
Figure 2. Prompt template for motivating examples.
XR Interaction. There are various ways in which users can interact with GUI elements. Some examples include: (1) Pressing buttons. Users can interact with GUI elements by pressing buttons on their devices, such as trigger buttons or joysticks on controllers. (2) Touching. Users can touch GUI elements by moving their devices to the location of elements in the virtual world. Some apps extend users’ hands in the virtual world with a special object to allow users to interact with elements remotely by making the special object contact with others. (3) Gazing. Users can interact with GUI elements by directing their gaze to them. Some headsets support eye tracking that enables interaction using only eye movements. (4) Gestures and Movements. Users can interact with GUI elements through gestures or movements, such as raising or waving hands. As more methods for human-computer interaction are proposed, the interaction methods are becoming more diverse.
@@ 2. Background on Extended Reality (XR)
Figure 3. LMM gives deviated IGE detection results on VR scenes (web, 2024g, d) because of aforementioned challenges
2.1. Motivating Examples
In this section, we present our motivational study to further demonstrate the technical challenges listed in section 1. Two authors randomly sample 30 VR apps from the popular Steam VR app store (web, 2023c), select screenshots from them, and use the prompt in Figure 2 to instruct LMM GPT-4o (web, 2024c) to identify and locate IGEs in the images.
As shown in figure 3(a), the mailbox is wrongly recognized as “Missile-like Object”. It is because the mailbox is viewed from an unusual angle, resulting in a special shape in the view and making it harder for the LMMs to identify semantics. This misunderstanding of GUI elements reveals the LMM’s inability to face the heterogeneous GUI elements in VR apps. For another example, figure 3(b), the bounding boxes in the image either cover only part of the object, locate objects mismatched with the label, or even mark the background. Although the LMM identifies most objects in the image, it fails to locate them well, revealing its insufficient spatial perception.
To summarize, the LMM alone cannot effectively detect IGEs in highly complex and varied XR apps. To boost the LMM on IGE detection in XR apps, it is necessary to provide the LMM with more context information and integrate the model with different techniques. Following this idea, we propose Orienter to imitate human behavior to enhance the LMM on detecting IGEs in XR apps.